What is the y-intercept of the function f(x) = –x + ?


1
Expert's answer
2021-11-15T18:28:42-0500

In mathematics, the y-intercept of a function, f(x), is the y-value of the point where the graph of f(x)crosses the y

y-axis. This occurs where x=0, so we can find the y-intercept of a function, f(x), by plugging x=0 into the function, which is equivalent to calculating f(0).

Where f(x) =-x+7

"\therefore" f(0)=-0+7

f(0)=7

y-intercept of f(x) =-x+7, is equal to 7

"\therefore" y-intercept =7